# Thresholded Lasso for high dimensional variable selection

### v2

@@ 2 The Thresholded Lasso estimator
Theorem 2.1 states that sparse oracle inequalities as elaborated in Section 1.1 hold for the Thresholded Lasso under no restriction on $\beta_{\min}$. Theorem 2.1 is the key contribution of this paper and is proved in the supplementary Section A.1. We do not optimize constants in this paper.
@@
(Ideal model selection for the Thresholded Lasso) Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $Y=X\beta+\epsilon$, where $\epsilon=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ is a vector containing independent and identically distributed (i.i.d.) noise with $\epsilon_{i}\sim N(0,\sigma^{2})$ for all $i\in[n]$. Suppose the columns of $X$ are normalized to have $\ell_{2}$ norm $\sqrt{n}$. Suppose $\textnormal{{RE}}(s_{0},4,X)$ holds with $K(s_{0},4)$, for $s_{0}$ as in (13), and the sparse eigenvalue conditions (5) and (6) hold. Let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\sqrt{2\log p/n}\sigma$ $\geq 2\lambda_{\sigma,a,p}$, where $d_{0}\geq 2\sqrt{1+a}$ for $a\geq 0$. Set $t_{0}=C_{4}\lambda\sigma\geq 2\sqrt{1+a}\lambda\sigma$, for some constant $C_{4}\geq D_{1}$ for $D_{1}$ as in (26). Let $D^{\prime}_{0}$ be as in (2.1). Set $I=\left\{j\in[p]:\beta_{j,\text{\rm init}}\geq t_{0}\right\}$. Set $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}_{I^{c}}=0$. Then with probability at least $1-{\mathbb{P}}\left({\mathcal{T}}_{a}^{c}\right)$, we obtain
@@
$\lVert\hat{\beta}-\beta\rVert_{2}^{2}\leq D_{4}^{2}s_{0}\lambda^{2}\sigma^{2},\text{ where}\;\;$ (17)
@@
Theorem 2.1 relies on new oracle results that we prove for the Lasso estimator under the same $\textnormal{{RE}}(s_{0},4,X)$ condition in Theorem 2.4. The Lasso estimator achieves essentially the same bound in terms of $\ell_{2}$ loss as stated in (11), which adapts nearly ideally not only to the uncertainty in the support set $S$ but also the “significant” set. It is clear from our analysis in [47] that, showing an oracle inequality as in (11) for the initial Lasso estimator $\beta_{\text{\rm init}}$, cf. Theorem 2.4, as well as applying new techniques for analyzing algorithms involving thresholding followed by OLS refitting will be crucial in proving the sparse oracle inequalities for the Thresholded Lasso. We discuss the initial estimator $\beta_{\text{\rm init}}$ and set $I$ in Section 2.1.
@@
Moreover, (17) implies that (11) holds for the Thresholded Lasso estimator. To see (14), we have by definition of $s_{0}$, where $0\leq s_{0}\leq s<p$,
@@
[…]
@@ 2.1 The thresholding rules
Consider the linear regression model in (1). Suppose we aim to target the set of variables of size at least $\sigma\sqrt{2\log p/n}$. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values, for $s_{0}$ as in (13). Lemma 2.3 is a deterministic result.
@@
(A deterministic result.) Let $\beta_{\text{\rm init}}$ be an initial estimator of a $s$-sparse $\beta$ in (1), where $\epsilon\sim N_{n}(0,\sigma^{2}I)$ and $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n},j\in[p]$. Let $\beta_{T_{0}}\in\mathbb{R}^{p}$ be the restriction of $\beta$ to the set $T_{0}$, where $T_{0}$ denotes the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$. Suppose for $\lambda:=\sqrt{{2\log p}/{n}}$,
@@
Then, a tighter bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}:=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}$ or $\lVert\beta_{\text{\rm init},T_{0}^{c}}\rVert_{2}$ will decrease the threshold $t_{0}$ while a tighter bound on $\lVert h_{T_{0}}\rVert_{2}:=\lVert(\beta-\beta_{\text{\rm init}})_{T_{0}}\rVert_{2}\leq D_{0}^{\prime}\lambda\sigma\sqrt{s_{0}}$ will tighten the upper bound in (21) on the bias component through the triangle inequality. In general, we allow $t_{0}$ to be chosen from a reasonably wide range, where we tradeoff the width of the range with the tightness of the upper bound on the $\ell_{2}$ loss for $\hat{\beta}$. This saves us from having to estimate incoherence parameters in a refined (and tedious) manner.
@@
(Oracle inequalities of the Lasso) Let $Y=X\beta+\epsilon$ for $\epsilon$ containing independent and identically distributed (i.i.d.) noise with $\epsilon_{i}\sim N(0,\sigma^{2})$ for all $i\in[n]$, and $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n}$ for all $j\in[p]$. Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $s_{0}$ be as in (13) and $T_{0}$ denote locations of the $s_{0}$ largest coefficients of $\beta$ in absolute values. Suppose $\textnormal{{RE}}(s_{0},4,X)$ holds with $K(s_{0},4)$ and the upper sparse eigenvalue condition (5) holds. Let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$, where $a\geq 0$ and $d_{0}\geq 2\sqrt{1+a}$. Let $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ be as in Lemma 2.3. Then on ${\mathcal{T}}_{a}$ as in (16),
@@
$\left\lVert h_{T_{0}}\right\rVert_{1}+\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\left\lVert h_{T_{0}}\right\rVert_{1}+\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}\leq D_{2}\lambda\sigma s_{0},\quad\text{and}$
@@
where $D_{0},\ldots,D_{3}$ are defined in (2.1) to (28). Moreover, for any subset $I_{0}\subset S$, by assuming that $\textnormal{{RE}}(|I_{0}|,4,X)$ holds with $K(|I_{0}|,4)$, we have
$\left\lVert X\beta_{\text{\rm init}}-X\beta\right\rVert_{2}/\sqrt{n}\leq\left\lVert X\beta-X\beta_{I_{0}}\right\rVert_{2}/\sqrt{n}+\frac{3}{2}K(|I_{0}|,4)\lambda_{n}\sqrt{|I_{0}|},$ (22)
where $\beta_{I_{0}}\in\mathbb{R}^{p}$ is the restriction of $\beta$ to the set $I_{0}$.
Theorem 2.4 may be of independent interest. We give a proof sketch in Section 3.1. The full proof of Theorem 2.4 appears in the supplementary Section H, which yields the following: cf. (19), (29), and (35),
$D_{0}$ $=$ $\left\{D\vee\sqrt{2}\big{(}d_{0}K^{2}(s_{0},4)+K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+2d_{0}K^{2}(s_{0},3)\big{)}\right\},$
$\asymp$ $\left\{D,\sqrt{2}\big{(}K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+d_{0}(2K^{2}(s_{0},3)+K^{2}(s_{0},4))\big{)}\right\},$
$D^{\prime}_{0}$ $=$ $\left\{D\vee[d_{0}K^{2}(s_{0},4)]\vee[K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}(s_{0},3)]\right\},$
[…]
$\ell(s_{0})=({\theta_{s_{0},2s_{0}}}/{\sqrt{\Lambda_{\min}(2s_{0})}})\wedge\sqrt{\Lambda_{\max}(s_{0})}$
$D_{1}$ $=$ $d_{0}\left\{[\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+\frac{9}{4}K^{2}(s_{0},3)]\vee 4K(s_{0},4)^{2}\vee\frac{3\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}\right\},$ (26)
$D_{2}$ $=$ $d_{0}\left\{[\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+4K^{2}(s_{0},3)]\vee 5K(s_{0},4)^{2}\vee\frac{4\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}\right\},$ (27)
$D_{3}$ $=$ $\sqrt{\Lambda_{\max}(s-s_{0})}+d_{0}K(s_{0},4)/2+d_{0}K(s_{0},3).$ (28)
@@
$\leq$ $5\big{(}\Lambda_{\max}(s-s_{0})/d_{0}\vee(d_{0}K(s_{0},4)^{2})\big{)}$
We can obtain an upper bound on $\theta_{s_{0},2s_{0}}$ in two ways. Let disjoint sets $J,J^{\prime}\subset[p]$ satisfy $\left\lvert J\right\rvert\leq s_{0}$ and $\left\lvert J^{\prime}\right\rvert\leq 2s_{0}$ and vectors $\upsilon,\upsilon^{\prime}$ satisfy $\left\lVert\upsilon\right\rVert_{2}=\left\lVert\upsilon^{\prime}\right\rVert_{2}=1$. We have by the Cauchy–Schwarz inequality and the parallelogram identity:
@@
We compare it with a well known $\ell_{p}$ error result by [3] (cf. Theorem 7.2) in Section 3.1. The sparse oracle properties of the Thresholded Lasso in terms of variable selection, $\ell_{2}$ loss, and prediction error then follow from Theorem 2.4, Lemmas 2.3 and 2.7; cf. Section A.1.
To help build intuition, we first state in Lemma 2.5 a general result on the $\ell_{2}$ loss for the OLS estimator when a subset of relevant variables is missing from the fixed model $I$. Lemma 2.5 is also an important technical contribution of this paper, which may be of independent interest. The assumption on $I$ being fixed is then relaxed in Lemma 2.7. We note that Lemmas 2.5 and 2.7 apply to $X$ so long as sparse eigenvalue conditions (5) and (6) hold.
@@
(OLS estimator with missing variables) Suppose sparse eigenvalue conditions (5) and (6) hold. Given a deterministic set $I\subset[p]$, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S=S\setminus I$. Let $\left|I\right|=m\leq(c_{0}s_{0})\wedge s$ for some absolute constant $c_{0}$. Suppose $|I\cup S|\leq 2s$. Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ be the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$. Then, with probability at least $1-2\exp(-3m/64)$,
@@
As a consequence of (6) and (5), for any subset $I$ such that $|I|\leq 2s$,
$\infty>\Lambda_{\max}(2s)$$\geq$ $\Lambda_{\max}(|I|)\geq\Lambda_{\max}\left({X_{I}^{T}X_{I}}/{n}\right)$
@@
Moreover, for disjoint sets $I$ and $S_{{\mathcal{D}}}=S\setminus I$ as in Lemmas 2.5 and 2.7, we have
@@
where it is understood that $\Lambda_{\min}(2s)=0$ is also permitted; cf. Proof of Lemma 2.8 in the supplementary Section E.
@@
[47] Suppose (6) and (5) hold. Given an arbitrary set $I\subset[p]$, possibly random, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S$. Suppose on event ${\mathcal{T}}_{a}$, we have $\left|I\right|=:m\leq(c_{0}s_{0})\wedge s$ for some absolute constant $c_{0}$ and $|I\cup S|\leq 2s$. Then, for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, it holds on ${\mathcal{T}}_{a}$ that
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$$\leq$ $\big{(}\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}+1\big{)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)},$ (32)
where $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ is the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$.
The proofs for Lemmas 2.5 and 2.7 are deferred to Section B. We now state Lemma 2.8, which follows from [8] (Lemma 1.2).
@@
[[8]] Suppose sparse eigenvalue conditions (5) and (6) hold. The following statements hold: (a) For all disjoint sets $I,S_{{\mathcal{D}}}\subseteq[p]$ of cardinality $|S_{{\mathcal{D}}}|<s$ and $|I|+|S_{{\mathcal{D}}}|\leq 2s$,
@@
(b) Without the lower sparse eigenvalue condition (6), following the same arguments leading to (29), we still obtain an upper bound under (5):
$\theta_{|I|,|S_{{\mathcal{D}}}|}\leq\Lambda_{\max}(s)\wedge(\Lambda_{\max}(2s)/2)\;\text{ if }\;\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s.$ (33)
@@
(I.) In Lemma 2.7, we have $\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s$; Moreover, the lower sparse eigenvalue condition (6) can be replaced with the following relaxed condition:

### v3

@@ 2 The Thresholded Lasso estimator
Theorem 2.1 states that sparse oracle inequalities as elaborated in Section 1.1 hold for the Thresholded Lasso under no restriction on $\beta_{\min}$. We do not optimize constants in this paper. Theorem 2.1 is the key contribution of this paper and is proved in Section 6.1.
@@
(Ideal model selection for the Thresholded Lasso) Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $s_{0}$ be as in (13). Let $Y=X\beta+\epsilon$, where $\epsilon=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ is a vector containing independent and identically distributed (i.i.d.) noise with $\epsilon_{i}\sim N(0,\sigma^{2})$ for all $i\in[n]$. Suppose the columns of $X$ are normalized to have $\ell_{2}$ norm $\sqrt{n}$. Suppose $\textnormal{{RE}}(s_{0},4,X)$ holds with $K(s_{0},4)$, and the sparse eigenvalue conditions (5) and (6) hold. Let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\sqrt{2\log p/n}\sigma$ $\geq 2\lambda_{\sigma,a,p}$, where $d_{0}\geq 2\sqrt{1+a}$ for $a\geq 0$. Set $t_{0}=C_{4}\lambda\sigma\geq 2\sqrt{1+a}\lambda\sigma$, for some constant $C_{4}\geq D_{1}$ for $D_{1}$ as in (27). Set $I=\left\{j\in[p]:\left\lvert\beta_{j,\text{\rm init}}\right\rvert\geq t_{0}\right\}$. Set $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}_{I^{c}}=0$. Then with probability at least $1-{\mathbb{P}}\left({\mathcal{T}}_{a}^{c}\right)$, we obtain
@@
[…]
@@
Theorem 2.1 relies on new oracle results that we prove for the Lasso estimator under $\textnormal{{RE}}(s_{0},4,X)$ and the upper sparse eigenvalue condition (5) in Theorem 2.4. The Lasso estimator achieves essentially the same bound in terms of $\ell_{2}$ loss as stated in (11), which adapts nearly ideally not only to the uncertainty in the support set $S$ but also to the size of the “significant” set; cf. Section 3.1. It is clear from our analysis in [47] that, showing an oracle inequality as in (11) for the initial Lasso estimator $\beta_{\text{\rm init}}$, cf. Theorem 2.4, as well as applying new techniques for analyzing algorithms involving thresholding followed by OLS refitting will be crucial in proving sparse oracle inequalities for the Thresholded Lasso. We discuss the initial estimator $\beta_{\text{\rm init}}$ and the set $I$ in Section 2.1. Moreover, (17) implies that (11) holds for the Thresholded Lasso estimator.
@@
To see (14), we have by definition of $s_{0}$, where $0\leq s_{0}\leq s<p$,
@@
$\text{ and}\;\;s_{0}\lambda^{2}\sigma^{2}$ $\geq$ $\sum_{j=1}^{s_{0}+1}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})\geq(s_{0}+1)\min(\beta_{s_{0}+1}^{2},\lambda^{2}\sigma^{2}),$
@@ 2.1 The thresholding rules
Consider the linear regression model in (1). Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values, for $s_{0}$ as in (13). Suppose we aim to select the set of variables of size at least $\sigma\sqrt{2\log p/n}$ while controlling the bias. Lemma 2.3 is a deterministic result toward this goal. We prove Lemma 2.3 in Section 6.2.
@@
(A deterministic result.) Let $\beta_{\text{\rm init}}$ be an initial estimator of the $s$-sparse $\beta$ in the linear model (1), where $\epsilon\sim N_{n}(0,\sigma^{2}I)$ and $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n},j\in[p]$. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $\beta_{T_{0}}$ be the restriction of $\beta$ to the set $T_{0}$. Let $\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})\in\mathbb{R}^{p}$ denote its 0-extended version such that $\beta_{T_{0}^{c}}^{\mathop{\natural}\kern 0.60275pt}({T_{0}})=0$ and $\beta_{T_{0}}^{\mathop{\natural}\kern 0.60275pt}({T_{0}})=\beta_{T_{0}}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Suppose for $\lambda:=\sqrt{{2\log p}/{n}}$,
@@
Then, a tighter bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}:=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}$ or $\lVert\beta_{\text{\rm init},T_{0}^{c}}\rVert_{2}$ will decrease the threshold $t_{0}$ while a tighter bound on $\lVert h_{T_{0}}\rVert_{2}:=\lVert(\beta-\beta_{\text{\rm init}})_{T_{0}}\rVert_{2}\leq D_{0}^{\prime}\lambda\sigma\sqrt{s_{0}}$ will tighten the upper bound in (21) on the bias component through the triangle inequality. In general, we allow $t_{0}$ to be chosen from a reasonably wide range, where we tradeoff the width of the range with the tightness of the upper bound on the $\ell_{2}$ loss for $\hat{\beta}$ in Step 3. This saves us from having to estimate incoherence parameters in a refined (and tedious) manner. Theorem 2.4 may be of independent interest. We give a proof sketch in Section 3.1.
@@
(Oracle inequalities of the Lasso) Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $s_{0}$ be as in (13). Suppose $\textnormal{{RE}}(s_{0},4,X)$ and the upper sparse eigenvalue condition (5) hold. Under the settings of Lemma 2.3, let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$, where $a\geq 0$ and $d_{0}\geq 2\sqrt{1+a}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ be defined as in Lemma 2.3. Then on ${\mathcal{T}}_{a}$,
@@
$\left\lVert h_{T_{0}}\right\rVert_{1}+\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\left\lVert\beta_{T_{0}}-\beta_{\text{\rm init},T_{0}}\right\rVert_{1}+\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}\leq D_{2}\lambda\sigma s_{0},\quad\text{and}$
@@
where $D_{0},\ldots,D_{3}$ are defined in (23) to (29). Moreover, for any subset $I_{0}\subset S$, by assuming that $\textnormal{{RE}}(|I_{0}|,4,X)$ holds with $K(|I_{0}|,4)$, we have
$\left\lVert X\beta_{\text{\rm init}}-X\beta\right\rVert_{2}/\sqrt{n}\leq\left\lVert X\beta-X_{I_{0}}\beta_{I_{0}}\right\rVert_{2}/\sqrt{n}+\frac{3}{2}K(|I_{0}|,4)\lambda_{n}\sqrt{|I_{0}|}.$ (22)
We compare Theorem 2.4 with the $\ell_{p}$ error bounds by [3] (cf. Theorem 7.2) in Section 3.1. The sparse oracle properties of the Thresholded Lasso in terms of variable selection, $\ell_{2}$ loss, and prediction error then follow from Theorem 2.4, Lemmas 2.3 and 2.7; cf. Section 6.1. The full proof of Theorem 2.4 appears in the supplementary Section D, which yields the following: cf. (19),
$D_{0}$ $=$ $D\vee\sqrt{2}\big(d_{0}K^{2}(s_{0},4)+K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+2d_{0}K^{2}(s_{0},3)\big),$ (23)
$D^{\prime}_{0}$ $=$ $D\vee(d_{0}K^{2}(s_{0},4))\vee(K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}(s_{0},3)),$ (26)
$\text{ where }\;D=\sqrt{\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}}\big(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\big)\;\text{ and }$
$\ell(s_{0})=({\theta_{s_{0},2s_{0}}}/{\sqrt{\Lambda_{\min}(2s_{0})}})\wedge\sqrt{\Lambda_{\max}(s_{0})},$
$D_{1}$ $=$ $d_{0}[\big(\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+\frac{9}{4}K^{2}(s_{0},3)\big)\vee 4K(s_{0},4)^{2}\vee\frac{3\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}],$ (27)
$D_{2}$ $=$ $d_{0}[\big(\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+4K^{2}(s_{0},3)\big)\vee 5K(s_{0},4)^{2}\vee\frac{4\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}],$ (29)
$\text{ and }\quad D_{3}=\sqrt{\Lambda_{\max}(s-s_{0})}+d_{0}K(s_{0},4)/2+d_{0}K(s_{0},3).$
@@
$\leq$ $5\big(\Lambda_{\max}(s-s_{0})/d_{0}\vee(d_{0}K(s_{0},4)^{2})\big).$
The bounds in (30) and (36) ensure that $\theta_{s_{0},2s_{0}}<\infty$ under (5) and $1/\Lambda_{\min}(2s_{0})<2K^{2}(s_{0},1)\leq 2K^{2}(s_{0},4)<\infty$ under $\textnormal{{RE}}(s_{0},4,X)$. In fact, we can obtain an upper bound on $\theta_{s_{0},2s_{0}}$ in two ways. Let disjoint sets $J,J^{\prime}\subset[p]$ satisfy $\left\lvert J\right\rvert\leq s_{0}$ and $\left\lvert J^{\prime}\right\rvert\leq 2s_{0}$ and vectors $\upsilon,\upsilon^{\prime}$ satisfy $\left\lVert\upsilon\right\rVert_{2}=\left\lVert\upsilon^{\prime}\right\rVert_{2}=1$. We have by the Cauchy–Schwarz inequality and the parallelogram identity:
@@
To help build intuition, we first state in Lemma 2.5 a general result on the $\ell_{2}$ loss for the OLS estimator, when a subset of relevant variables is missing from the fixed model $I$. Lemma 2.5 is also an important technical contribution of this paper, which may be of independent interest. The assumption on $I$ being fixed is then relaxed in Lemma 2.7. Given $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, let $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ be the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$ in both lemmas. Let $c_{0},c>0$ be some absolute constants.
@@
(OLS estimator with missing variables) Suppose sparse eigenvalue conditions (5) and (6) hold. Given a deterministic set $I\subset[p]$, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S=S\setminus I$. Let $\left|I\right|=m\leq(c_{0}s_{0})\wedge s$. Suppose $|I\cup S|\leq 2s$. Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$. Then, with probability at least $1-2\exp(-3m/64)$,
@@
Lemmas 2.5 and 2.7 apply to $X$ so long as sparse eigenvalue conditions (5) and (6) hold. As a consequence, for any subset $I$ such that $|I|\leq 2s$,
$\infty>\Lambda_{\max}(2s)$ $\geq$ $\Lambda_{\max}(|I|)\geq\Lambda_{\max}\left({X_{I}^{T}X_{I}}/{n}\right)$ (31)
@@
For disjoint sets $I$ and $S_{{\mathcal{D}}}=S\setminus I$ as in Lemmas 2.5 and 2.7, we have
@@
where it is understood that $\Lambda_{\min}(2s)=0$ is also permitted; cf. Lemma 2.8 and its proof in the supplementary Section B.
@@
[47] Suppose sparse eigenvalue conditions (5) and (6) hold. Given an arbitrary set $I\subset[p]$, possibly random, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S$. Suppose on event ${\mathcal{T}}_{a}$, $\left|I\right|\leq(c_{0}s_{0})\wedge s$ and $|I\cup S|\leq 2s$. Then, for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ as defined in Lemma 2.5, it holds on ${\mathcal{T}}_{a}$ that
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $\leq$ $\big(\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}+1\big)\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)}.$ (33)
The proofs for Lemmas 2.5 and 2.7 are deferred to Sections 6.4 and 6.5. We now state Lemma 2.8, which follows from [8].
@@
[Lemma 1.2 [8]] Suppose sparse eigenvalue conditions (5) and (6) hold. The following statements hold: (a) For all disjoint sets $I,S_{{\mathcal{D}}}\subseteq[p]$ of cardinality $|S_{{\mathcal{D}}}|<s$ and $|I|+|S_{{\mathcal{D}}}|\leq 2s$,
@@
(b) Without the lower sparse eigenvalue (LSE) condition (6), we still obtain
$\theta_{|I|,|S_{{\mathcal{D}}}|}\leq\Lambda_{\max}(s)\wedge(\Lambda_{\max}(2s)/2)<\infty\;\text{ if }\;\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s,$ (34)
under the USE condition (5), following the same arguments leading to (30).
@@
(I.) In Lemma 2.7, we have $\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s$. Hence, the LSE condition (6) can be replaced with the following relaxed lower sparse eigenvalue condition:
@@
2.2 Discussions
The tight analysis in Theorem 2.1 for the Thresholded Lasso estimator is motivated by the sparse oracle inequalities on the Gauss-Dantzig selector under the UUP, which is originally shown to hold in a conference paper by [47]. Previously, it has been shown that (11) holds with high probability for the Dantzig selector under the condition of a Uniform Uncertainty Principle (UUP), where the UUP states that for all $s$-sparse sets $J$, the columns of $X$ corresponding to $J$ are almost orthogonal [9].